OpenStudy (anonymous):

SOMEONE HELP PLEASE. What is the point of inflection of the following equation ( see reply ).

OpenStudy (anonymous):

\[e ^{-2x}\]

OpenStudy (anonymous):

Take the derivative twice and then set it to zero

OpenStudy (anonymous):

i know. could you help me do that?

OpenStudy (anonymous):

dy/dx=-2e^(-2x) take the derivative again: 4e^-2x

OpenStudy (anonymous):

ok i got the second derivative i don't know how to solve for zero with that though.

OpenStudy (anonymous):

2nd derivative = 4e^-2x. Set it equal to zero to find any values you can use to set up a second derivative test.

OpenStudy (anonymous):

O wait I see what you mean.. so 4e^-2x=0 e^-2x=0 -2x=ln0 -2x=-infiniti x=infiniti

OpenStudy (anonymous):

There only exists an inflection point when the graph shifts from concave up and concave down.

OpenStudy (anonymous):

so the graph doesn't shift from concave up or down?.. basically there is no point of inflection? because in my textbook there is an answer.

OpenStudy (anonymous):

What is the answer in your text book.

OpenStudy (anonymous):

in fact there are two points of inflection in my textbook.

OpenStudy (anonymous):

lol i have no clue then :)

OpenStudy (anonymous):

haha ok thanks anyways.

OpenStudy (anonymous):

one of the points is (-1,root2)

OpenStudy (anonymous):

i don't have the textbook with me though so i'm not 100 percent sure thats right.

OpenStudy (anonymous):

if you have a graphing calculator on you, try to graph that equation. it might just be the case that you wrote the problem down wrong. I don't know though. Inflection points only occur if the graph shifts from concave up and concave down and vice versa. I cant seem to find any values to even attempt a second derivative test for inflection points so...

OpenStudy (anonymous):

ok, thanks for your help.